OpenStudy (anonymous):

Factor completely. 3r^2-3 If a polynomial is prime, state this

OpenStudy (callisto):

1. take out the common factor 3r^2-3 = 3 (r^2 -1) 2. use the identity a^2 - b^2 = (a+b)(a-b) to factorise r^2 -1 = 3(r-1)(r+1)

OpenStudy (anonymous):

Is that the final answer for factoring the equation?

OpenStudy (callisto):

What do you think ?

OpenStudy (anonymous):

It looks factored to me. Thanks.

OpenStudy (callisto):

Welcome :)